# IRConStyle: Image Restoration Framework Using Contrastive Learning and Style Transfer

### 24 Feb 2024 v1

@@ Abstract
Recently, the contrastive learning paradigm has achieved remarkable success in high-level tasks such as classification, detection, and segmentation. However, contrastive learning applied in low-level tasks, like image restoration, is limited, and its effectiveness is uncertain. This raises a question: Why does the contrastive learning paradigm not yield satisfactory results in image restoration? In this paper, we conduct in-depth analyses and propose three guidelines to address the above question. In addition, inspired by style transfer and based on contrastive learning, we propose a novel module for image restoration called ConStyle, which can be efficiently integrated into any U-Net structure network. By leveraging the flexibility of ConStyle, we develop a general restoration network for image restoration. ConStyle and the general restoration network together form an image restoration framework, namely IRConStyle. To demonstrate the capability and compatibility of ConStyle, we replace the general restoration network with transformer-based, CNN-based, and MLP-based networks, respectively. We perform extensive experiments on various image restoration tasks, including denoising, deblurring, deraining, and dehazing. The results on 19 benchmarks demonstrate that ConStyle can be integrated with any U-Net-based network and significantly enhance performance. For instance, ConStyle NAFNet significantly outperforms the original NAFNet on SOTS outdoor (dehazing) and Rain100H (deraining) datasets, with PSNR improvements of 4.16 dB and 3.58 dB with 85% fewer parameters. The code is available at: https://github.com/Dongqi-Fan/IRConStyle
@@ 4 Experiments
ConStyle Restormer 34.33 31.71 28.51 34.89 32.66 29.64 34.37 31.73 28.52 35.01 32.72 29.70
Differ. +0.01 0 0 +0.02 +0.01 +0.01 0 -0.01 -0.01 +0.01 -0.03 -0.01
@@
ConStyle NAFNet 34.31 31.69 28.50 34.82 32.58 29.55 34.34 31.71 28.51 34.91 32.64 29.58
Differ. +0.05 +0.05 +0.04 +0.20 +0.22 +0.26 +0.02 0 -0.01 +0.11 -0.01 -0.05
@@
ConStyle MAXIM-1S 30.54/0.9051 33.80/0.9424 26.54/0.8244 29.22/0.8930
Differ. +0.04/ +0.0012 -0.09/ -0.0007 +0.06/ +0.0014 +0/ +0.0012
@@ 4.3 Experiment Results
ConStyle MAXIM-1S 30.44/0.9710 29.20/0.8845 36.23/0.9748
Differ. -0.03/ -0.0026 -0.30/ -0.0027 +0.69/ +0.0031

### 07 Mar 2024 v3

@@ Abstract
Recently, the contrastive learning paradigm has achieved remarkable success in high-level tasks such as classification, detection, and segmentation. However, in low-level tasks, like image restoration, contrastive learning is limited, and its effectiveness is uncertain. This raises a question: Why does the contrastive learning paradigm not yield remarkable results in image restoration? In this paper, we conduct in-depth analyses and propose three guidelines to address the above question. In addition, inspired by style transfer and based on contrastive learning, we propose a novel module for image restoration called ConStyle, which can be efficiently integrated into any U-Net structure network. By leveraging the flexibility of ConStyle, we develop a general restoration network for image restoration. The ConStyle and general restoration network together form an image restoration framework, namely IRConStyle. To demonstrate the capability and compatibility of ConStyle, we replace the general restoration network with transformer-based, CNN-based, and MLP-based networks, respectively. We perform extensive experiments on various image restoration tasks, including denoising, deblurring, deraining, and dehazing. The results on 19 benchmarks demonstrate that ConStyle can be integrated with any U-Net-based network and significantly enhance the performance of the network. For instance, ConStyle NAFNet significantly outperforms the original NAFNet on SOTS outdoor (dehazing) and Rain100H (deraining) datasets, with PSNR improvements of 4.16 dB and 3.58 dB with 85% fewer parameters. The code is available at: https://github.com/Dongqi-Fan/IRConStyle
@@ 4 Experiments
ConStyle Restormer 34.33 31.71 28.51 34.89 32.66 29.64 34.37 31.74 28.52 35.01 32.74 29.71
Differ. +0.01 0 0 +0.02 +0.01 +0.01 0 +0 +0 -0.01 -0.01 +0
@@
ConStyle NAFNet 34.31 31.69 28.50 34.82 32.58 29.55 34.34 31.71 28.52 34.91 32.64 29.63
Differ. +0.05 +0.05 +0.04 +0.20 +0.22 +0.26 +0.02 0 +0 +0.11 -0.01 +0
@@
ConStyle MAXIM-1S 30.77/0.9093 33.93/0.9435 26.54/0.8261 29.26/0.8951
Differ. +0.27/ +0.0054 +0.04/ +0.0004 +0.06/ +0.0031 +0.04/ +0.0033
@@ 4.3 Experiment Results
ConStyle MAXIM-1S 30.59/0.9713 29.20/0.8845 36.23/0.9748
Differ. +0.12/ -0.0023 -0.30/ -0.0027 +0.69/ +0.0031